# Sequential Monte-Carlo testing by betting

### v4

@@ Abstract
In a Monte-Carlo test, the observed dataset is fixed, and several resampled or permuted versions of the dataset are generated in order to test a null hypothesis that the original dataset is exchangeable with the resampled/permuted ones. Sequential Monte-Carlo tests aim to save computational resources by generating these additional datasets sequentially one by one, and potentially stopping early. While earlier tests yield valid inference at a particular prespecified stopping rule, our work develops a new anytime-valid Monte-Carlo test that can be continuously monitored, yielding a p-value or e-value at any stopping time possibly not specified in advance. Despite the added flexibility, it significantly outperforms the well-known method by Besag and Clifford, stopping earlier under both the null and the alternative without compromising power. The core technical advance is the development of new test martingales (nonnegative martingales with initial value one) for testing exchangeability against a very particular alternative. These test martingales are constructed using new and simple betting strategies that smartly bet on the relative ranks of generated test statistics. The betting strategies are guided by the derivation of a simple log-optimal betting strategy, have closed form expressions for the wealth process, provable guarantees on resampling risk, and display excellent power in practice.
@@ 1 Introduction
Suppose we have some data $X_{0}$, from which we calculate a test statistic $T(X_{0})$. Suppose further that we have the ability to generate additional data $X_{1},X_{2},\dots,X_{B}$ and calculate test statistics $T(X_{1}),T(X_{2}),\dots,T(X_{B})$. Denoting $Y_{i}=T(X_{i})$, assume further that $Y_{1},\dots,Y_{B}$ are always exchangeable conditional on $Y_{0}$. Suppose that we are trying to test the null hypothesis
$H_{0}:Y_{0},Y_{1},\dots Y_{B}\text{ are exchangeable.}$.
@@
$\mathfrak{p}^{\mathrm{perm}}_{B}:=\frac{1+\sum_{b=1}^{B}\mathbf{1}\{Y_{b}\geq Y_{0}\}}{B+1}.$
Use the shorthand $Y_{0}^{B}:=\{Y_{0},\dots,Y_{B}\}$, it will be useful to denote the numerator above by
$\mathrm{rank}(Y_{0};Y_{0}^{B}):=\sum_{b=0}^{B}\mathbf{1}\{Y_{b}\geq Y_{0}\}.$
But how should one choose $B$? In practice, it is often just set as a suitably large constant like 1000, or sometimes even larger if multiple testing corrections need to be performed. The goal of this paper is to describe a simple “anytime-valid p-value” and an “e-process” for $H_{0}$ that is obtained using the principle of testing by betting.
Before we do this, it makes sense to first describe the implicit alternative:
$H_{1}:Y_{1},\dots,Y_{B}\text{ are exchangeable conditional on }Y_{0},\text{ but }Y_{0}\text{ is stochastically larger than }Y_{1},\dots,Y_{B}.$
While there do exist sequential tests for exchangeability, they have focused on (for example) changepoint [30] or Markov alternatives [20, 25]. The technique introduced in this paper is different (and in some ways simpler), and focuses on the above $H_{1}$.
@@
In order to avoid ties, each test statistic $Y_{i}$ should be considered as coming together with a random sample $\theta_{i}$ from $U[0,1]$ which is independent of all test statistics and all other random samples [32, 30]. If two test statistics are equal $Y_{i}=Y_{j}$, their random samples decide which test statistic is considered as “greater”. Hence, the event $\{Y_{b}\geq Y_{0}\}$ should be read as $\{Y_{b}>Y_{0}\}\cup\{Y_{b}=Y_{0},\theta_{b}>\theta_{0}\}$. To the detriment of power, drawing random samples could be avoided when a conservative choice can be made in case of a tie. For instance, if we count all ties between $Y_{0}$ and a generated test statistics $Y_{b}$ as $Y_{b}\geq Y_{0}$, the permutation p-value $\mathfrak{p}^{\mathrm{perm}}$ would still be valid, but conservative. The same holds for our binomial mixture strategy with uniform prior presented in Section 5.2.
@@
We can generate new datasets $X_{1}=(X,W_{1}),\ldots,X_{B}=(X,W_{B})$ by randomly permuting $W_{0}$ $B$ times. Note that we only permute the treatment labels and not the observed response data $X$. In this way, the datasets $X_{0},\ldots,X_{B}$ are exchangeable under the null hypothesis, but $X_{0}$ is not exchangeable with $X_{1},\ldots,X_{B}$ if the null hypothesis is not true. In addition, the generated datasets $X_{1},\ldots,X_{B}$ are always exchangeable (even conditional on $X_{0}$), since they were generated in an identical manner. The same properties hold for $Y_{0},Y_{1},\ldots,Y_{B}$, where $Y_{i}=T(X_{i})$ and $T$ can be any test statistic — a common choice is the difference in mean between the treated and untreated observations. Hence, this example fits perfectly into our introduced setting, and the reader should keep it in mind in what follows. We also consider it in our simulations (Section 7.17.3) and a real data application (Section 7.4).
@@
For example, $W_{0}$ could be a binary treatment vector, $X$ the observed responses and $Z$ contains the patient covariates. If the distribution of $W_{0}|Z$ is known, one can test $H_{0}^{\mathrm{CI}}$ using a conditional randomization test (CRT); while this is a common idea in causal inference, Candès et al. [5] discuss broader applications. For example, $W_{0}|Z$ is indeed known if we are analyzing data from a designed experiment where the treatment was randomly assigned to each patient based on a known probability that depends on that patient’s covariates. Here, we cannot permute $W_{0}$ because the subjects are not exchangeable conditional on $Z$, but we can redraw $W_{0}$. Let $W_{1},\ldots,W_{B}$ be sampled independently from the distribution of $W_{0}|Z$ and $Y_{b}:=T(W_{i},X,Z)$, $b\in\{0,1,\ldots,B\}$, for an arbitrary test statistic $T$. The test statistics $Y_{1},\ldots,Y_{B}$ are always exchangeable conditional on $Y_{0}$, while $Y_{0},Y_{1},\ldots,Y_{B}$ are exchangeable under $H_{0}^{\mathrm{CI}}$, matching the setting of our paper. The CRT p-value essentially uses the same formula as the permutation p-value $\mathfrak{p}^{\mathrm{perm}}$ on $Y_{0},\ldots,Y_{B}$ (even though it is technically not a permutation test, but a randomization test). However, generating samples from $W_{0}|Z$ can be computationally challenging or recalculating the test statistic on these samples can be very costly, so a lot of computational effort can potentially be saved by using our sequential permutation tests. We revisit this example in Section 7.5.
@@ 1.1 Our approach at a high level
We design a game of chance, where a gambler begins with one dollar, makes bets in each round so that their wealth changes over time. The value of the wealth at any point will be a measure of evidence against the null.
The game will proceed in rounds, which we index by $t=1,2,\dots$ — in round $t$, the gambler will first bet on the rank of $Y_{t}$ amongst $Y_{0},\dots,Y_{t}$, denoted $R_{t}=\mathrm{rank}(Y_{t};Y_{0}^{t})$. Then nature will draw $X_{t}$ and calculate $Y_{t}$ and reveal $R_{t}$ (hiding $Y_{t}$ itself), thus determining our payoff for that round.
It is important that we do not know the raw $Y$ values, but only the ranks $R$ are revealed in each round. So after the first round, we only know if $Y_{1}$ was larger or smaller than $Y_{0}$, but not the values of either $Y_{0}$ or $Y_{1}$.
The gambler’s bet will be specified by a nonnegative $t+1$-dimensional vector, denoted $B_{t}$. $B_{t}(r)$ denotes the amount their wealth gets multiplied by when $R_{t}=r$. Their wealth after $B$ rounds of betting is
$W_{B}:=\prod_{t=1}^{B}B_{t}(R_{t}).$
Our bet will have to satisfy the constraint $\mathbb{E}_{H_{0}}[B_{t}(R_{t})\mid R_{1}^{t-1}]=1$ so that the wealth process is a test martingale (a nonnegative martingale with initial value one). Under the null, the rank $R_{t}=\mathrm{rank}(Y_{t};Y_{0}^{t})$ is independent of $R_{1},\ldots,R_{t-1}$ and discretely uniformly distributed with equal mass at each $r=1,\dots,t+1$ [22], so the constraint simplifies to
$\sum_{r=1}^{t+1}B_{t}(r)=t+1.$
@@
for any stopping time $\tau$ with respect to the filtration $(\mathcal{R}_{t})_{t\geq 1}$ of the ranks $\mathcal{R}_{t}:=\sigma(R_{1},\dots,R_{t})$. Said differently, $W_{\tau}$ is an e-value at any stopping time $\tau$ that is possibly not even specified in advance. E-processes capture evidence against the null: the larger the wealth, the more the evidence against the null. If the null is true, (1) implies that no betting strategy and stopping rule is guaranteed to make money (i.e., demonstrate evidence against the null). However, if the alternative is true, then a good betting strategy can make money and the evidence against the null can grow over time (as more permutations are drawn). We will design such good betting strategies in this paper.
@@
$\mathfrak{p}_{t}:=1/(\sup_{s\leq t}W_{s})$
is an anytime-valid p-value, or “p-process” [14, 13, 21], meaning that
$\mathbb{P}_{H_{0}}(\mathfrak{p}_{\tau}\leq\alpha)\leq\alpha,$
at any arbitrary stopping time $\tau$.
@@
Optional Input: Stopping rule $\mathcal{S}$, potentially data-dependent and decided on the fly.
@@
2:for $t=1,2,...$ do
3:     Choose betting strategy $B_{t}=(B_{t}(1),\ldots,B_{t}(t+1))$ with $\sum_{r=1}^{t+1}B_{t}(r)=t+1$
4:     Reveal $R_{t}$
5:     $W_{t}=W_{t-1}\cdot B_{t}(R_{t})$
6:     if $\mathcal{S}(R_{1},\ldots,R_{t})=\text{stop}$ then
7:         $\tau=t$
8:         return $\tau,W_{\tau},(\max_{s=1,\ldots,\tau}W_{s})^{-1}$
9:     end if
10:end for
@@ 1.2 Related work
The use of test martingales and “testing by betting” have been popularized as a safe approach to sequential, anytime-valid inference [27, 26, 13, 12, 36]. An overview of this new methodology and the recent advances is given in [21].
@@
There have been several other works on sequential tests of exchangeability [30, 20, 25, 16]. The most important differentiating factor of the current paper from those is the alternative hypothesis being tested. The aforementioned works involve the data coming in sequentially, and the alternative involves either (a) a changepoint from exchangeability (i.e., the first $k$ samples are exchangeable, but not the following samples, with $k=0$ being a special case), (b) Markovian alternatives (i.e., there is some serial time-ordered dependence between the data). Our paper considers the setting where the data are fixed, the permutations are drawn sequentially, and thus — by design — the alternative is such that the first test statistic is special while the rest are exchangeable. In some sense, our paper is really targeted at a sequential version of the classical problem of permutation testing, while the earlier papers are essentially trying test if the data are i.i.d., or if there is some distribution shift or drift. Despite their different goals, we describe some key technical aspects of these papers in Appendix A. Another related martingale-based approach by Waudby-Smith and Ramdas [35] constructs confidence sets for some parameter by sampling without replacement (WoR). However, this problem implicitly tests a different null hypothesis, which leads to a dissimilar method.
@@
The closest related work to ours is the famous sequential Monte-Carlo p-values method of Besag and Clifford [3], who introduced a simple strategy for reducing the number of permutations drawn that is still very popular today. It can be shown that for a certain parameter choice, their method leads to the same decisions as the classical permutation test while needing less permutations [28]. Inspired by this, further sequential Monte-Carlo tests with the objective to reduce the number of permutations were proposed [9, 29]. An important difference to our approach is that theirs are stopped based on fixed predefined rules, and thus yield valid p-values only at these particular stopping times, and are not valid at any other stopping time. In addition, the existing approaches do not have the option to continue testing by drawing more permutations once the stopping criterium is reached (for example, if the null is not rejected), while our method can always handle drawing additional permutations and stopping at a later point, even if these flexibilities are not specified in advance. Finally, we will experimentally compare our approach to the Besag-Clifford method. Our simulations suggest that our method can stop sooner than theirs under both the null and the alternative, without sacrificing power.

### v5

@@ Abstract
In a Monte-Carlo test, the observed dataset is fixed, and several resampled or permuted versions of the dataset are generated in order to test a null hypothesis that the original dataset is exchangeable with the resampled/permuted ones. Sequential Monte-Carlo tests aim to save computational resources by generating these additional datasets sequentially one by one, and potentially stopping early. While earlier tests yield valid inference at a particular prespecified stopping rule, our work develops a new anytime-valid Monte-Carlo test that can be continuously monitored, yielding a p-value or e-value at any stopping time possibly not specified in advance. It generalizes the well-known method by Besag and Clifford, allowing it to stop at any time, but also encompasses new sequential Monte-Carlo tests that tend to stop sooner under the null and alternative without compromising power. The core technical advance is the development of new test martingales (nonnegative martingales with initial value one) for testing exchangeability against a very particular alternative. These test martingales are constructed using new and simple betting strategies that smartly bet on whether a generated test statistic is greater or smaller than the observed one. The betting strategies are guided by the derivation of a simple log-optimal betting strategy, have closed form expressions for the wealth process, provable guarantees on resampling risk, and display excellent power in practice.
@@ 1 Introduction
Suppose we have some data $X_{0}$, from which we calculate a test statistic $S(X_{0})$. Suppose further that we have the ability to generate additional data $X_{1},X_{2},\dots$ and calculate test statistics $S(X_{1}),S(X_{2}),\dots$ Denoting $Y_{i}=S(X_{i})$, assume further that $Y_{1},Y_{2},\dots$ are always exchangeable conditional on $Y_{0}$. Suppose that we are trying to test the null hypothesis
$H_{0}:Y_{0},Y_{1},\dots\text{ are exchangeable.}$.
@@
$Q^{\mathrm{perm}}_{T}:=\frac{1+\sum_{t=1}^{T}\mathbbm{1}\{Y_{t}\geq Y_{0}\}}{T+1}$ (1)
for some prespecified number of permutations $T$. But how should one choose $T$? In practice, it is often just set as a suitably large constant like 1000, or sometimes even larger if multiple testing corrections need to be performed. The goal of this paper is to describe a simple “anytime-valid p-value” (or “p-process”) and an “e-process” for $H_{0}$ that is obtained using the principle of testing by betting.
Before we do this, it makes sense to first describe the (working) alternative:
$H_{1}:Y_{1},Y_{2},\dots\text{ are exchangeable conditional on }Y_{0},\text{ but }Y_{0}\text{ is stochastically larger than }Y_{1},Y_{2},\dots\text{ }.$
We used the word “working” above to highlight that when the null is false, we will not require or assume that the above $H_{1}$ is true. Since the complement of $H_{0}$ is enormous, no test can be powerful against all alternatives, and specifying such a “working” alternative helps guide the design of powerful and practical tests.
While there do exist sequential tests for exchangeability, they have focused on (for example) changepoint [36] or Markov alternatives [25, 30]. The technique introduced in this paper is different, and focuses on the above $H_{1}$ that is practically relevant for settings like two-sample or independence testing using permutations.
@@
In order to avoid ties, each test statistic $Y_{i}$ should be considered as coming together with a random sample $\theta_{i}$ from $U[0,1]$ which is independent of all test statistics and all other random samples [38, 39, 36]. If two test statistics are equal $Y_{i}=Y_{j}$, then comparing $\theta_{i},\theta_{j}$ decides which test statistic is considered as “greater”. Hence, the event $\{Y_{t}\geq Y_{0}\}$ should be read as $\{Y_{t}>Y_{0}\}\cup\{Y_{t}=Y_{0},\theta_{t}>\theta_{0}\}$. To the detriment of power, drawing random samples can be avoided when a conservative choice can be made in case of a tie. For instance, if we count all ties between $Y_{0}$ and a generated test statistics $Y_{t}$ as $Y_{t}\geq Y_{0}$, the permutation p-value $Q^{\mathrm{perm}}$ would still be valid, but conservative. The same holds for our binomial mixture strategy with uniform density presented in Section 5.2.
Remark 2.
In a permutation test one can either draw the permutations without replacement or with replacement. If it is computationally feasible to draw all possible permutations, one usually does the former and sets the parameter $T$ in (1) to equal the total number of possible permutations. However, it is computationally more convenient to draw the permutations with replacement. In the hypothesis formulation above it is implicitly assumed that we have access to an infinite number of generated test statistics and therefore that the permutations are drawn with replacement. This allows us to characterize the asymptotic behavior in the remainder of the paper. However, our proposed methods are also valid if permutations are sampled without replacement and/or a maximum number of permutations is fixed in advance.
@@
We can generate new datasets $X_{1}=(X,W_{1}),X_{2}=(X,W_{2}),\ldots$ by randomly permuting $W_{0}$. Note that we only permute the treatment labels and not the observed response data $X$. In this way, the datasets $X_{0},X_{1},\ldots$ are exchangeable under the null hypothesis, but $X_{0}$ is not exchangeable with $X_{1},X_{2},\ldots$ if the null hypothesis is not true. In addition, the generated datasets $X_{1},X_{2},\ldots$ are always exchangeable (even conditional on $X_{0}$), since they were generated in an identical manner. The same properties hold for $Y_{0},Y_{1},\ldots$, where $Y_{i}=S(X_{i})$ and $S$ can be any test statistic — a common choice is the difference in mean between the treated and untreated observations. Hence, this example fits perfectly into our introduced setting, and the reader may keep it in mind in what follows. We also consider it in our simulations (Section 7.17.3) and a real data application (Section 7.4).
@@
For example, $W_{0}$ could be a binary treatment vector, $X$ the observed responses and $Z$ contains the patient covariates. If the distribution of $W_{0}|Z$ is known, one can test $H_{0}^{\mathrm{CI}}$ using a conditional randomization test (CRT); while this is a common idea in causal inference, Candès et al. [5] discuss broader applications. For example, $W_{0}|Z$ is indeed known if we are analyzing data from a designed experiment where the treatment was randomly assigned to each patient based on a known probability that depends on that patient’s covariates. Here, we cannot permute $W_{0}$ because the subjects are not exchangeable conditional on $Z$, but we can redraw $W_{0}$ as follows. Let $W_{1},W_{2},\ldots$ be sampled independently from the distribution of $W_{0}|Z$ and $Y_{t}:=S(W_{t},X,Z)$, $t\in\mathbb{N}_{0}$, for an arbitrary test statistic $S$. The test statistics $Y_{1},Y_{2},\ldots$ are always exchangeable conditional on $Y_{0}$, while $Y_{0},Y_{1},\ldots$ are exchangeable under $H_{0}^{\mathrm{CI}}$, matching the setting of our paper. The CRT p-value essentially uses the same formula as the permutation p-value (1) (even though it is technically not a permutation test, but a randomization test). Generating samples from $W_{0}|Z$ can be computationally challenging or recalculating the test statistic on these samples can be very costly, so a lot of computational effort can potentially be saved by using our sequential permutation tests. We revisit this example in Section 7.5.
@@ 1.1 Our approach at a high level
We design a game of chance, where a gambler begins with one dollar, makes bets in each round so that their wealth changes over time. The value of the wealth at any point will be a measure of evidence against the null. Throughout the paper, $\mathbb{P}$ denotes the true but unknown probability distribution and $\mathbb{E}$ the corresponding expected value, whereas $\mathbb{P}_{H_{0}}$ and $\mathbb{E}_{H_{0}}$ denote the probability and expectation under the null hypothesis (i.e. corresponding to an arbitrary distribution that satisfies $H_{0}$).
The game will proceed in rounds, which we index by $t=1,2,\dots$ — in round $t$, the gambler will first bet on the outcome of $I_{t}:=\mathbbm{1}\{Y_{t}\geq Y_{0}\}$. We will soon specify what properties this bet must satisfy. Then nature will draw $X_{t}$ and calculate $Y_{t}$ and reveal $I_{t}$ (hiding $Y_{t}$ itself), thus determining our payoff for that round. In the following we will also refer to the event $\{I_{t}=1\}$ as a “loss” and to $\{I_{t}=0\}$ as a “win”. This is because under the alternative, it is intuitive that the gambler will usually put more money on the event $\{I_{t}=0\}$, but we also theoretically justify this in Section 3. Similarly, the random variable $L_{T}:=\sum_{t=1}^{T}I_{t}$ will be called the number of losses after $T$ permutations.
It is important that we do not know the raw $Y$ values, but only the indicators $I$ are revealed in each round. So after the first round, we only know if $Y_{1}$ was larger or smaller than $Y_{0}$, but not the values of either $Y_{0}$ or $Y_{1}$.
The gambler’s bet will be specified by a betting function $B_{t}:\{0,1\}\rightarrow\mathbb{R}_{\geq 0}$. $B_{t}(r)$ denotes the amount their wealth gets multiplied by when $I_{t}=r$, $r\in\{0,1\}$. Their wealth after $T$ rounds of betting is
$W_{T}:=\prod_{t=1}^{T}B_{t}(I_{t}).$
Denoting $I_{1}^{t}=\{I_{1},\dots,I_{t}\}$, our bet will have to satisfy the constraint $\mathbb{E}_{H_{0}}[B_{t}(I_{t})\mid I_{1}^{t-1}]=1$ so that the wealth process is a test martingale (a nonnegative martingale with initial value one under $H_{0}$). Of course, this implies that only one of $B_{t}(0)$ and $B_{t}(1)$ can be larger than 1, and typically this will be $B_{t}(0)$ (as guided by $H_{1}$). To simplify the above constraint, note that conditional on $I_{1}^{t-1}$, the event $\{I_{t}=1\}$ is equivalent to the event that $Y_{t}$ ranks among the $L_{t-1}+1$ largest test statistics of $Y_{0},\ldots,Y_{t}$. Since $Y_{0},\ldots,Y_{t}$ are exchangeable under $H_{0}$, the constraint simplifies to
$B_{t}(0)\frac{t-L_{t-1}}{t+1}+B_{t}(1)\frac{1+L_{t-1}}{t+1}=1.$ (2)
@@
for any stopping time $\tau$ with respect to the filtration $(\mathcal{I}_{t})_{t\geq 1}$ of the indicators $\mathcal{I}_{t}:=\sigma(I_{1},\dots,I_{t})$. Recall that an integer-valued random variable $\tau$ is called stopping time, if $\{\tau=t\}$ is measurable with respect to $\mathcal{I}_{t}$ for all $t\in\mathbb{N}$.
Said differently, (3) states that $W_{\tau}$ is an e-value at any stopping time $\tau$ that is possibly not even specified in advance. E-processes capture evidence against the null: the larger the wealth, the more the evidence against the null. If the null is true, (3) implies that no betting strategy and stopping rule is guaranteed to make money (i.e., demonstrate evidence against the null). However, if the alternative is true, then a good betting strategy can make money and the evidence against the null can grow over time (as more permutations are drawn). We will design such good betting strategies in this paper.
@@
$Q_{t}:=1/(\sup_{s\leq t}W_{s})$ (5)
is an anytime-valid p-value, or “p-process” [15, 14, 26], meaning that at any arbitrary stopping time $\tau$, $Q_{\tau}$ is a p-value:
$\mathbb{P}_{H_{0}}(Q_{\tau}\leq\alpha)\leq\alpha.$
We notate the p-values with the capital letter $Q$ and their realizations with the lowercase letter $q$ to distinguish them from the parameters $p$ introduced later.
@@
Optional Input: Stopping rule $\mathcal{S}$ (potentially data-dependent and decided on the fly).
@@
2:$L_{0}=0$
3:for $t=1,2,...$ do
4:     Choose a bet $B_{t}=(B_{t}(0),B_{t}(1))$ with $B_{t}(0)\frac{t-L_{t-1}}{t+1}+B_{t}(1)\frac{1+L_{t-1}}{t+1}=1$
5:     Reveal $I_{t}$
6:     Calculate $L_{t}=L_{t-1}+I_{t}$
7:     $W_{t}=W_{t-1}\cdot B_{t}(I_{t})$
8:     if $\mathcal{S}(I_{1},\ldots,I_{t})=\text{stop}$ then
9:         $\tau=t$
10:         return $\tau,W_{\tau},(\max_{s=1,\ldots,\tau}W_{s})^{-1}$
11:     end if
12:end for
Remark 3.
A general approach for the sequential testing of exchangeability has been proposed by Vovk and colleagues [39, 38, 36] based on conformal prediction. Instead of betting at each step $t$ on the indicator $I_{t}=\mathbbm{1}\{Y_{t}\geq Y_{0}\}$, Vovk considers the rank $R_{t}=\sum_{i=0}^{t}\mathbbm{1}\{Y_{i}\geq Y_{t}\}$ of $Y_{t}$ among $Y_{0},\ldots,Y_{t}$. We could have proceeded in the same way. However, since $Y_{1},Y_{2},\ldots$ are exchangeable conditional on $Y_{0}$ under our alternative, the rank of $Y_{t}$ among $Y_{1},\ldots,Y_{t}$ is uniform (even conditional on $R_{1},\ldots,R_{t-1}$). Therefore, the only relevant information is whether $Y_{t}\geq Y_{0}$. Another subtle difference to Vovk’s approach is the randomization technique used to handle ties (see Remark 2). In Vovk’s approach, one has to draw a random sample $\theta_{t}$ for every test statistic $Y_{t}$, while we only require a random sample $\theta_{t}$ if $Y_{t}$ and $Y_{0}$ actually tie. This reduces the computational effort, the main goal of this paper. In addition, with Vovk’s randomization approach one must specify a bet for each value in $[0,1]$, while we only bet on $\{0,1\}$ which simplifies the betting function $B_{t}$ and leads to closed forms for the wealth $W_{T}$, as we will illustrate in the remainder of the paper.
To summarize, our approach can be considered as a special case of Vovk’s conformal martingale approach. However, our amendments simplify the method substantially for our particular alternative — an alternative, which to the best of our knowledge has not been stated in the literature before and naturally arises from the classical problem of sequential permutation testing. The development of powerful tests for this alternative requires the construction of new betting strategies.
@@ 1.2 Related work
The use of test martingales and “testing by betting” have been popularized as a safe approach to sequential, anytime-valid inference [32, 31, 14, 13, 43]. An overview of this new methodology and the recent advances is given by Ramdas et al. [26].
@@
There have been several other works on sequential tests of exchangeability [36, 25, 30, 17, 20]. The most important differentiating factor of the current paper from those is the alternative hypothesis being tested. The aforementioned works involve the data coming in sequentially, and the alternative involves either (a) a changepoint from exchangeability (i.e., the first $j$ samples are exchangeable, but not the following samples, with $j=0$ being a special case), (b) Markov alternatives (i.e., there is some serial time-ordered dependence between the data).
In contrast, our paper considers the setting where the data are fixed, the permutations are drawn sequentially, and thus — by design — the alternative is such that the first test statistic is special while the rest are exchangeable.
In some sense, our paper is really targeted at a sequential version of the classical problem of permutation testing, while the earlier papers are essentially trying to test if the data are i.i.d., or if there is some distribution shift or drift. Despite their different goals, we describe some key technical aspects of these papers in Appendix A. A more detailed description of the relation to the conformal prediction approach by Vovk and colleagues [38, 39, 36] is given in Remark 3. Another related martingale-based approach by Waudby-Smith and Ramdas [42] constructs confidence sets for some parameter by sampling without replacement (WoR). However, this problem implicitly tests a different null hypothesis, which leads to a dissimilar method.
@@
The closest related work to ours is the famous sequential Monte-Carlo p-values method of Besag and Clifford [3], who introduced a simple strategy for reducing the number of permutations drawn that is still very popular today. It can be shown that for a certain parameter choice, their method leads to the same decisions as the classical permutation test while needing less permutations [33]. Inspired by this, further sequential Monte-Carlo tests with the objective to reduce the number of permutations were proposed [10, 34]. An important difference to our approach is that theirs are stopped based on fixed predefined rules, and thus yield valid p-values only at these particular stopping times, and are not valid at any other stopping time. In addition, the existing approaches do not have the option to continue testing by drawing more permutations once the stopping criterium is reached (for example, if the null is not rejected), while our method can always handle drawing additional permutations and stopping at a later point, even if these flexibilities are not specified in advance.
Our general betting approach generalizes the Besag-Clifford method, but also allows to construct new strategies. Our simulations suggest that these new strategies can stop sooner under both the null and the alternative, without sacrificing power.
There are also other approaches to reduce the computational cost of permutation tests than sequential permutation p-values. Koning [18] and Koning and Hemerik [19] consider gaining computational efficiency by considering a subgroup of permutations. Another line of work tries to approximate the null distribution of $Y_{0}$ analytically [27, 7, 23]. Gandy [11] proposes to bound the probability of obtaining a different decision than the limiting permutation p-value. A recent and comprehensive overview of the literature on computational efficient permutation tests — including an assessment of the approach introduced in this paper — is given by Stoepker and Castro [35].